# Fairness Evaluation of Large Language Models in Academic Library Reference Services

### v2

@@ Abstract
As libraries explore large language models (LLMs) for use in virtual reference services, a key question arises: Can LLMs serve all users equitably, regardless of demographics or social status? While they offer great potential for scalable support, LLMs may also reproduce societal biases embedded in their training data, risking the integrity of libraries’ commitment to equitable service. To address this concern, we evaluate whether LLMs differentiate responses across user identities by prompting six state-of-the-art LLMs to assist patrons differing in sex, race/ethnicity, and institutional role. We found no evidence of differentiation by race or ethnicity, and only minor evidence of stereotypical bias against women in one model. LLMs demonstrated nuanced accommodation of institutional roles through the use of linguistic choices related to formality, politeness, and domain-specific vocabularies, reflecting professional norms rather than discriminatory treatment. These findings suggest that current LLMs show a promising degree of readiness to support equitable and contextually appropriate communication in academic library reference services.
@@
1 Introduction
Large language models (LLMs) represent a transformative opportunity for academic libraries to democratize access to information services (Cox and Tzoc,, 2023; Wang et al.,, 2024; Mahajan,, 2025). They can provide responses to patron inquiries regardless of physical constraints such as time, location, and staffing. In addition, cognitive barriers posed by lower education levels (Klomsri and Tedre,, 2016; Yevelson-Shorsher and Bronstein,, 2018) or language differences (Sin and Kim,, 2018; Zhao et al.,, 2021) will no longer limit service accessibility. Imagine a library patron being able to access research assistance at 2 AM, or a student receiving personalized database recommendations during finals week, when human librarians are overwhelmed with requests. However, realizing this potential requires careful attention to equity concerns that have long challenged human reference services. Audit studies have documented disparities in virtual and email reference interactions, where patrons are treated differently based on perceived race or gender (Shachaf and Horowitz,, 2006; Vladoiu et al.,, 2023). These patterns underscore persistent concerns about implicit bias in reference services. The urgency of this evaluation has been emphasized in a recent statement from the International Federation of Library Associations and Institutions (IFLA)’s “Entry Point to Libraries and AI” framework, which identifies bias as a risk requiring immediate attention before AI deployment in libraries (IFLA,, 2025).
@@ 1 Introduction
In this work, we conducted a fairness evaluation using synthetic patron personas across six state-of-the-art LLMs, including both open models and proprietary, commercial giants. The evaluation procedure, termed the Fairness Evaluation Protocol (FEP), combines automated bias detection with linguistic analysis to assess whether LLMs systematically vary their responses based on patron demographics or institutional roles. It characterizes general behavioral patterns and model-specific tendencies, and supports interpretation of LLM behavior through a sociolinguistic lens.
@@
Our analysis yields three main findings. First, LLMs are insensitive to racial and ethnic cues. Second, LLMs generally treat men and women equitably, with only one model evincing minor stereotypical gender associations. Third, LLMs demonstrate sophisticated institutional role accommodation through strategic use of formality strategies and domain-specific vocabulary, reflecting a nuanced understanding of professional communication rather than discriminatory treatment.
@@
These findings demonstrate the potential of today’s LLMs to provide equitable service to patrons with diverse demographic characteristics and to deliver contextually appropriate professional communication in academic library reference services. Given the variation across models in communication style, libraries must carefully select and prompt LLMs to align with their desired communication norms. Finally, we contribute a generalizable, interpretable, and model-agnostic fairness evaluation protocol that enables libraries and similar institutions to audit the fairness and understand the behavior of their LLM-based services.
2 Literature Review
2.1 Equitable Service in Academic Libraries & Related Challenges
@@
The American Library Association (ALA) Code of Ethics begins with the principle: “We provide the highest level of service to all library users through appropriate and usefully organized resources; equitable service policies; equitable access; and accurate, unbiased, and courteous responses to all requests” (American Library Association,, 2021). However, these professional commitments to equitable service face persistent challenges in practice.
@@
Human reference interactions, despite librarians’ best intentions, have shown documented disparities based on patron demographics. Shachaf and Horowitz, (2006) conducted a seminal audit study revealing that virtual reference queries signed with African American or Arab-sounding names received responses that were less complete and less courteous than those signed with Caucasian-sounding names. This finding was later replicated in UK libraries, where Hamer, (2021) found that email inquiries from “Black African” personas were less likely to receive helpful responses than those from “White British” counterparts. More recently, Vladoiu et al., (2023) examined reference interactions and found that a patron profile with an African name received the least helpful service, while a profile with an East Asian name received the most comprehensive assistance. In addition to demographic disparities, variation by institutional role has emerged as another axis of concern, with Shachaf and Horowitz, (2006) documenting that faculty and enrolled students often receive more attentive service than alumni, staff, or unaffiliated users.
@@
Algorithmic bias in information systems has emerged as a parallel concern for librarians. Noble, (2018) demonstrated how search engines can reinforce racism through biased results, while Reidsma, (2019) brought these concerns directly to library discovery systems, documenting how discovery tools can return systematically different and potentially biased results for equivalent queries. These findings illustrate that libraries have long grappled with bias across both human and technological dimensions of service delivery.
@@
As libraries increasingly explore LLM integration for reference services, these longstanding concerns about bias have become more pressing. IFLA’s manifestos on AI use acknowledge ethical concerns around bias, highlighting the need for libraries to understand and address these issues proactively before deployment (IFLA,, 2024, 2025). Recent research on AI in academic libraries similarly identifies algorithmic bias as “one of the primary concerns” that “can lead to unintended discrimination in search results and content recommendations” (Islam et al.,, 2025). This convergent professional concern highlights how the evaluation of LLM fairness represents not just a technical necessity, but a fundamental extension of libraries’ longstanding commitment to equitable service.
2.2 LLMs: Applications & Fairness Evaluation
@@
LLMs have demonstrated capabilities in engaging with patrons, showcasing their potential in fields such as healthcare, education, and industry. In healthcare, LLMs enhance diagnostics (Saab et al.,, 2025) and clinical decision support (Gaber et al.,, 2025; Schaye et al.,, 2025; Rajaganapathy et al.,, 2025). In travel planning, they demonstrate versatility in itinerary creation and personalization tasks (Ren et al.,, 2024). These implementations highlight LLMs’ capacity for complex, contextualized assistance, making their adoption in academic libraries both feasible and potentially impactful.
@@
This transformative potential, however, does not come without significant fairness concerns. LLMs are trained on vast text corpora that inevitably contain historical biases, which they can reproduce in their outputs (Dhamala et al.,, 2021). Bender et al., (2021) cautioned that ever-larger training sets can make LLMs “stochastic parrots” that regurgitate societal prejudices embedded in their source data. For example, the BOLD benchmark (Dhamala et al.,, 2021) found that open-ended generations from prominent LLMs exhibit more bias than human-authored Wikipedia text. The StereoSet benchmark, which aims to understand and mitigate stereotypical biases regarding gender, profession, race, and religion, reported that well-known language models such as BERT and GPT-2 exhibit strong stereotypical biases (Nadeem et al.,, 2021). Other benchmarks, such as CrowS-Pairs (Nangia et al.,, 2020) and WinoBias (Zhao et al.,, 2018), have revealed systematic biases across additional dimensions, such as age, disability, physical appearance, and sexual orientation.
@@
The LLM community has made substantial efforts to improve model safety and respect through techniques such as reinforcement learning from human feedback (RLHF) (Ouyang et al.,, 2022; Touvron et al.,, 2023) and constitutional AI methods (Bai et al.,, 2022). However, balancing helpfulness with safety remains challenging, and these methods typically target overall safety rather than domain-specific applications. Recent surveys (Gallegos et al.,, 2024; Chu et al.,, 2024) acknowledge improvements in newer models such as GPT-4 but emphasize that no model is entirely bias-free, especially in the flexible, generative settings exemplified by reference dialogue. Meta-analyses of LLM fairness research (Sheng et al.,, 2019) have highlighted the importance of domain-specific evaluation, as bias patterns can vary significantly across application contexts and user populations. For instance, Hofmann et al., (2024) demonstrated that LLMs harbor raciolinguistic stereotypes, offering less helpful or more critical responses when prompted with dialects associated with African American English, despite maintaining surface-level politeness. This finding suggests that bias can manifest through subtle linguistic patterns rather than overt discrimination.
@@
Fairness literature draws a useful conceptual distinction between group fairness and individual fairness. Group fairness refers to the notion that outcomes should not differ across groups defined by sensitive attributes (e.g., race or gender), while individual fairness suggests similar individuals should be treated similarly (Dwork et al.,, 2012; Mehrabi et al.,, 2021). Our evaluation protocol, as described in Section 4.3, primarily tests for group fairness: we assess whether model outputs vary systematically with demographic or institutional variables, holding prompt content constant. By examining telltale words, we can infer whether observed differences reflect bias or appropriate personalization, a distinction that speaks to the well-known fairness trade-off between consistency and contextualization (Barocas et al.,, 2019; Binns,, 2020). We return to this trade-off in the discussion (Section 6).
@@
Given their role as stewards of knowledge resources serving patrons from diverse backgrounds, academic libraries occupy a distinctive position in which fairness deserves in-depth investigation before LLMs are deployed. To this end, we propose a fairness evaluation method, detailed in Section 4.3, that is tailored to the library context (but generalizable to other applications; see Section 7). The first stage is to detect words in LLM outputs that are strongly associated with specific demographic groups or institutional status. These telltale words reveal differentiation among patrons of different demographic attributes or institutional roles.
2.3 Linguistic Theories for Analyzing Institutional Communication
@@
The words we detected span a spectrum from signals of discriminatory treatment, to stereotypical bias, to appropriate domain-specific customization. To analyze these linguistic patterns, we draw on several key theories that illuminate the functional, social, and contextual dimensions of language in this institutional setting.
@@
A foundational perspective is offered by Halliday,’s systemic functional linguistics framework (Halliday,, 1985), which posits that language is a resource for making meaning in social contexts. This theory identifies three core “metafunctions:” the ideational (conveying information), the interpersonal (managing social relationships), and the textual (organizing the message). Academic librarians must fulfill all three functions simultaneously: delivering accurate information, attending to patrons’ diverse social identities, and structuring communications appropriately for the institutional context. Consequently, language in this context, including that generated by an LLM, must effectively serve these three interconnected purposes.
@@
The communicative context of librarianship can be analyzed through Communication Accommodation Theory (CAT) (Giles et al.,, 1991). CAT suggests that individuals strategically and often unconsciously adapt their communication to manage social distance and project a desired identity. Recent developments in CAT (Gallois et al.,, 2005) show that accommodation strategies are moderated by cultural and social context, making our chosen setting particularly relevant for understanding these dynamics. Academic libraries represent an ideal context for examining accommodation theory, as they involve systematic interactions across clearly defined institutional hierarchies—from undergraduate students to faculty to external users—within a professional service framework. In this scenario, a librarian might adjust their language based on the perceived expertise, identity, and institutional role of the patron. These adjustments may involve convergence or divergence. Convergence involves mirroring the patron’s language to foster solidarity, while divergence involves using more formal language to signify professional boundaries. Ideally, a well-functioning LLM should also be capable of calibrating its tone across various patron types, yet in a way that avoids reinforcing inequitable social hierarchies.
@@
Central to managing these interactions is the concept of politeness. Politeness theory (Brown and Levinson,, 1987) highlights how interactions are strategically managed through face-saving behaviors. Linguistic strategies balance patrons’ needs for appreciation and inclusion (termed “positive face”) with their desire for autonomy and non-imposition (i.e., “negative face”). For instance, the utterance “I’d be happy to help you find that resource” mainly addresses positive politeness through offering help. It also indirectly addresses negative face because the utterance is phrased as voluntary, rather than imposing an obligation.
@@
Politeness use is not uniform across social identities, however. Research by Holmes, (1995) and Herring, (1994) shows that politeness strategies can vary significantly according to gender. For example, Herring, (1994) observed that women tend to use more positive politeness markers, including hedging, appreciation, and collaborative tone, while men more frequently engage in adversarial styles, such as direct disagreement or trolling behavior. However, more recent scholarship (Mills,, 2003) has challenged the notion that women are “necessarily always more polite than men,” emphasizing that politeness is highly context-dependent and that both women and men can be equally polite or impolite depending on the situation. Hence, it is important to evaluate LLMs within realistic professional contexts, as we do in this study, to understand whether models reproduce gendered or status-based biases in interactional tone.
@@
Finally, the language of librarianship operates within a distinct communicative genre shaped by the institutional context of academia (Swales,, 1990; Drew and Heritage,, 1992). This genre shapes expectations around tone, structure, and relational positioning, reinforcing norms of politeness, neutrality, and inclusiveness aligned with libraries’ professional values. A key dimension for analyzing this genre is formality. Drawing on register theory (Biber,, 1988) and theoretical frameworks by Irvine, (1979) and Heylighen and Dewaele, (1999), we operationalize formality as the degree to which language employs conventional professional markers such as formal address terms (“Dear” vs. “Hi” in greetings), structured closings (“Best Regards” vs. “Thanks” in sign-offs) and explicit rather than colloquial expressions. In institutional contexts, higher-status or more distant relationships (e.g., librarian to faculty or external users) typically elicit more formal, negative-politeness strategies that are respectful, indirect, and deferential. In contrast, interactions with peers or students may employ a more informal style with positive politeness markers. An equitable LLM is expected to modulate its formality to align with professional expectations for each patron type.
@@ 3 Research Questions
    How do current LLMs compare in delivering reference services, and what similarities and differences emerge in their behavioral patterns?
@@
The third question addresses LLMs’ pragmatic responses across different patron types within professional settings. In institutional contexts, communication strategies often reflect distinct social and institutional positions. In academic libraries specifically, patron type correlates with role, status, and resource access—distinctions that may signal broader variations in educational level, professional authority, or institutional privilege. Examining how LLMs respond to patrons based on their stated affiliation (e.g., faculty, student, staff, or unaffiliated user) allows us to assess whether AI systems encode assumptions or preferences that parallel social hierarchies. We adopt this patron-type taxonomy from Montana State University’s LibStats (Jordan,, 2008) reference question logs, a classification scheme representative of those used in academic library settings.
@@
4 Methodology
To answer our research questions, we simulated a scenario wherein an academic library user sends a common reference query via email and an LLM configured as a helpful, respectful, and honest librarian responds with a single message. In these queries, each user’s name is carefully synthesized to provide cues about their sex and race/ethnicity (see Section 4.1). Additionally, their patron type is explicitly indicated at the end of the email to make the LLM aware of their institutional role. This section describes our study design for simulating user-LLM interactions, the open and commercial LLMs under investigation and their configurations, and the framework we used to evaluate fairness in LLM-delivered service.
4.1 Synthesizing User Queries
We synthesize user queries as emails directed to virtual librarians at academic libraries (see an example in Section 4.1). Each query consists of three components: a query template, essentially boilerplate reflecting queries that academic librarians frequently encounter; a name, which identifies the email sender; and a patron type, which indicates their institutional role.
@@ 4.1 Synthesizing User Queries
We adopt three query templates based on real-world virtual reference interactions, following categories used in prior studies (Shachaf et al.,, 2008):
@@
Name & Patron Type
Names often carry implicit cues about an individual’s sex and race/ethnicity. When users sign emails with their names, these demographic signals may influence whether an LLM provides equitable service across different identity groups. We construct a balanced cohort of synthetic English names across twelve demographic groups, defined by all pairwise combinations of sex (male, female) and race/ethnicity (White; Black or African American; Asian or Pacific Islander; American Indian or Alaska Native; Two or More Races; Hispanic or Latino).
@@
We then sampled surnames from the U.S. Census Bureau’s 2010 surname dataset, which includes annotated distributions reflecting realistic racial/ethnic compositions. Each surname in the dataset is associated with a distribution over the aforementioned race/ethnicity categories. The sampling process first selects a surname uniformly at random, then samples a race/ethnicity label according to that surname’s normalized distribution. For example, even though the surname Wang is most commonly associated with Asian identity (95.2%), it is also recorded as 2.6% White and 0.3% Black in the dataset; in rare (yet realistic) cases, someone named John Wang might identify as White or Black. For race/ethnicity assignment, unlike sex assignment, we use a rejection-based method: we find a surname with a nonzero probability for the desired group and draw until the sampled label matches the target demographic.
This sampling process ensures that all twelve demographic groups are equally represented in the LLM interactions. Additionally, each synthetic identity was randomly assigned to one of six patron categories representing institutional roles: Alumni, Faculty, Graduate Student, Undergraduate Student, Staff, or Outside User.
@@
We plug the user’s name and patron type into a randomly chosen query template. An example assembled email reads:
@@
This populated query serves as the user prompt to the language model, analogous to a user’s interaction on the ChatGPT interface. A corresponding system prompt configures the LLM to act as a reference librarian from a randomly chosen institution (e.g., for the above prompt, the system prompt reads: “You are a helpful, respectful, and honest librarian from Louisiana State University.”). [^3] For models that support system prompts (all except Gemma-2), instructions are split into system and user messages. For Gemma-2, the system prompt is prepended to the user prompt. This ensures all LLMs are equivalently prompted.
4.2 Models & Experimental Setup
4.2.1 LLMs Under Study
We evaluate six state-of-the-art LLMs commonly used in research and industry. These fall into two categories based on their accessibility, licensing, and deployment models.

### v3

@@ Abstract
As libraries explore large language models (LLMs) for use in virtual reference services, a key question arises: Can LLMs serve all users equitably, regardless of demographics or social status? While they offer great potential for scalable support, LLMs may also reproduce societal biases embedded in their training data, risking the integrity of libraries’ commitment to equitable service. To address this concern, we evaluate whether LLMs differentiate responses across user identities by prompting six state-of-the-art LLMs to assist patrons differing in sex, race/ethnicity, and institutional role. We find no evidence of differentiation by race or ethnicity, and only minor evidence of stereotypical bias against women in one model. LLMs demonstrate nuanced accommodation of institutional roles through the use of linguistic choices related to formality, politeness, and domain-specific vocabularies, reflecting professional norms rather than discriminatory treatment. These findings suggest that current LLMs show a promising degree of readiness to support equitable and contextually appropriate communication in academic library reference services.
@@
Introduction
Large language models (LLMs) represent a transformative opportunity for academic libraries to democratize access to information services (cox2023chatgpt; wang2024science; mahajan2025democratization). They can provide responses to patron inquiries regardless of physical constraints such as time, location, and staffing. In addition, cognitive barriers posed by lower education levels (klomsri2016poor; yevelson2018three) or language differences (sin2018we; zhao2021information) need not limit service accessibility. Imagine a library patron being able to access research assistance at 2 AM, or a student receiving personalized database recommendations during finals week, when human librarians are overwhelmed with requests. However, realizing this potential requires careful attention to equity concerns that have long challenged human reference services. Audit studies have documented disparities in virtual and email reference interactions, where patrons are treated differently based on perceived race or gender (shachaf2006virtual; vladoiu2023gender). These patterns underscore persistent concerns about implicit bias in reference services. The urgency of this evaluation is emphasised in a recent statement from the International Federation of Library Associations and Institutions (IFLA)’s “Entry Point to Libraries and AI” framework, which identifies bias as a risk requiring immediate attention before AI deployment in libraries (ifla2025ai).
@@ Introduction
In this work, we conduct a fairness evaluation using synthetic patron personas across six state-of-the-art LLMs, including both open models and proprietary commercial giants. The evaluation procedure, termed the Fairness Evaluation Protocol (FEP), combines automated bias detection with linguistic analysis to assess whether LLMs systematically vary their responses based on patron demographics or institutional roles. It characterises general behavioural patterns and model-specific tendencies, and supports interpretation of LLM behaviour through a sociolinguistic lens.
@@
Our analysis yields three main findings. First, LLMs are insensitive to racial and ethnic cues. Second, LLMs generally treat men and women equitably, with only one model evincing minor stereotypical gender associations. Third, LLMs demonstrate sophisticated institutional role accommodation through calibrated use of formality and domain-specific vocabulary, reflecting a nuanced understanding of professional communication rather than discriminatory treatment.
@@
These findings demonstrate the potential of today’s LLMs to provide equitable service to patrons with diverse demographic characteristics and to deliver contextually appropriate professional communication in academic library reference services. Given the variation across models in communication style, libraries must carefully select and prompt LLMs to align with their desired communication norms. Finally, we contribute a generalizable, interpretable, and model-agnostic fairness evaluation protocol that enables libraries and similar institutions to audit the fairness and understand the behaviour of their LLM-based services.
Literature review
Equitable service in academic libraries and related challenges
@@
The American Library Association (ALA) Code of Ethics begins with the principle: “We provide the highest level of service to all library users through appropriate and usefully organized resources; equitable service policies; equitable access; and accurate, unbiased, and courteous responses to all requests” (ala2021ethics). However, these professional commitments to equitable service face persistent challenges in practice.
@@
Human reference interactions, despite librarians’ best intentions, have shown documented disparities based on patron demographics. shachaf2006virtual conducted a seminal audit study revealing that virtual reference queries signed with African American or Arab-sounding names received responses that were less complete and less courteous than those signed with Caucasian-sounding names. This finding was later replicated in UK libraries, where hamer2021colour found that email inquiries from “Black African” personas were less likely to receive helpful responses than those from “White British” counterparts. More recently, vladoiu2023gender examined reference interactions and found that a patron profile with an African name received the least helpful service, while a profile with an East Asian name received the most comprehensive assistance. In addition to demographic disparities, variation by institutional role has emerged as another axis of concern, with shachaf2006virtual documenting that faculty and enrolled students often receive more attentive service than alumni, staff, or unaffiliated users.
@@
Algorithmic bias in information systems has emerged as a parallel concern for librarians. noble2018algorithms demonstrated how search engines can reinforce racism through biased results, while reidsma2019masked brought these concerns directly to library discovery systems, documenting how discovery tools can return systematically different and potentially biased results for equivalent queries. These findings illustrate that libraries have long grappled with bias across both human and technological dimensions of service delivery.
@@
As libraries increasingly explore LLM integration for reference services, these longstanding concerns about bias have become more pressing. IFLA’s manifestos on AI use acknowledge ethical concerns around bias, highlighting the need for libraries to understand and address these issues proactively before deployment (ifla2024ai; ifla2025ai). Recent research on AI in academic libraries similarly identifies algorithmic bias as “one of the primary concerns” that “can lead to unintended discrimination in search results and content recommendations” (kumar2024artificial). This convergent professional concern highlights how the evaluation of LLM fairness represents not just a technical necessity, but a fundamental extension of libraries’ longstanding commitment to equitable service.
LLMs: applications and fairness evaluation
@@
LLMs have demonstrated capabilities in engaging with patrons, showcasing their potential in fields such as healthcare, education, and industry. In healthcare, LLMs enhance diagnostics (saab2025advancing) and clinical decision support (gaber2025evaluating; schaye2025large; rajaganapathy2025synoptic). In travel planning, they demonstrate versatility in itinerary creation and personalization tasks (ren2024large). These implementations highlight LLMs’ capacity for complex, contextualized assistance, making their adoption in academic libraries both feasible and potentially impactful.
@@
This transformative potential, however, does not come without significant fairness concerns. LLMs are trained on vast text corpora that inevitably contain historical biases, which they can reproduce in their outputs (dhamala2021bold). bender2021dangers cautioned that ever-larger training sets can make LLMs “stochastic parrots” that regurgitate societal prejudices embedded in their source data. For example, the BOLD benchmark (dhamala2021bold) found that open-ended generations from prominent LLMs exhibit more bias than human-authored Wikipedia text. The StereoSet benchmark, which aims to understand and mitigate stereotypical biases regarding gender, profession, race, and religion, reported that well-known language models such as BERT and GPT-2 exhibit strong stereotypical biases (nadeem2021stereoset). Other benchmarks, such as CrowS-Pairs (nangia2020crowspairs) and WinoBias (zhao2018gender), have revealed systematic biases across additional dimensions, such as age, disability, physical appearance, and sexual orientation.
@@
The LLM community has made substantial efforts to improve model safety and respect through techniques such as reinforcement learning from human feedback (RLHF) (ouyang2022training; touvron2023llama2) and constitutional AI methods (bai2022constitutional). However, balancing helpfulness with safety remains challenging, and these methods typically target overall safety rather than domain-specific applications. Recent surveys (gallegos2024bias; chu2024fairness) acknowledge improvements in newer models such as GPT-4 but emphasize that no model is entirely bias-free, especially in the flexible, generative settings exemplified by reference dialogue. Meta-analyses of LLM fairness research (sheng2019woman) have highlighted the importance of domain-specific evaluation, as bias patterns can vary significantly across application contexts and user populations. For instance, hofmann2024ai demonstrated that LLMs harbour raciolinguistic stereotypes, offering less helpful or more critical responses when prompted with dialects associated with African American English, despite maintaining surface-level politeness. This finding suggests that bias can manifest through subtle linguistic patterns rather than overt discrimination.
We ground our discussion of group fairness in the straightforward idea that model outputs should not systematically vary with protected attributes. In LLM evaluation, this serves as an initial check for differences across demographic categories, while recent theory cautions that fully satisfying this requirement alongside other aims can be impossible in some settings or may entail utility trade-offs (anthis2025impossibility). Applied studies adopt the same lens in varied contexts, for example when auditing decisions on tabular prompts (tayebi2025evaluating), monitoring educational AI for subgroup differences (chinta2024fairaied), and quantifying demographic skew in AI-generated co-authorship networks (kalhor2025measuring). In our setting, we operationalise this idea by testing whether the distribution of model responses is statistically indistinguishable across sex and race/ethnicity; only when deviations arise do we proceed to examine whether they reflect accommodation or bias.
@@
Fairness literature draws a useful conceptual distinction between group fairness and individual fairness. Group fairness refers to the notion that outcomes should not differ across groups defined by sensitive attributes (e.g., race or gender), while individual fairness requires that similar individuals, defined by a task-specific similarity metric, be treated similarly (dwork2012fairness; mehrabi2021survey). In this study, we operationalise similarity for reference service through institutional role: faculty members are similar to other faculty, undergraduates are similar to other undergraduates, and so forth. Individual fairness permits response variations across these institutional categories (e.g., providing more technical details for faculty versus making content more accessible to undergraduates), while requiring equivalent treatment within categories regardless of protected characteristics such as sex or race. Our evaluation protocol, as described in Fairness Evaluation Protocol section, primarily tests for group fairness: we assess whether model outputs vary systematically with demographic or institutional variables, holding prompt content constant. By examining telltale words, we can infer whether observed differences reflect bias or appropriate personalization, a distinction that speaks to the well-known fairness trade-off between consistency and contextualization (barocas2019fairness; binns2020apparent).
@@
Given their role as stewards of knowledge resources serving patrons from diverse backgrounds, academic libraries occupy a distinctive position in which fairness deserves in-depth investigation before LLMs are deployed. To this end, we propose a fairness evaluation method, detailed in Fairness Evaluation Protocol section, that is tailored to the library context (but generalizable to other applications; see Conclusion section). The first stage is to detect words in LLM outputs that are strongly associated with specific demographic groups or institutional status. These telltale words reveal differentiation among patrons of different demographic attributes or institutional roles.
Linguistic theories for analysing institutional communication
@@
The words we detected span a spectrum from signals of discriminatory treatment, to stereotypical bias, to appropriate domain-specific customization. To analyse these linguistic patterns, we draw on several key theories that illuminate the functional, social, and contextual dimensions of language in this institutional setting.
@@
A foundational perspective is offered by halliday1985introduction’s systemic functional linguistics framework (halliday1985introduction), which posits that language is a resource for making meaning in social contexts. This theory identifies three core “metafunctions:” the ideational (conveying information), the interpersonal (managing social relationships), and the textual (organising the message). Academic librarians must fulfil all three functions simultaneously: delivering accurate information, attending to patrons’ diverse social identities, and structuring communications appropriately for the institutional context. Consequently, language in this context, including that generated by an LLM, must effectively serve these three interconnected purposes.
@@
The communicative context of librarianship can be analysed through Communication Accommodation Theory (CAT) (giles1991accommodation). CAT suggests that individuals strategically and often unconsciously adapt their communication to manage social distance and project a desired identity. Recent developments in CAT (gallois2005communication) show that accommodation strategies are moderated by cultural and social context, making our chosen setting particularly relevant for understanding these dynamics. Academic libraries represent an ideal context for examining accommodation theory, as they involve systematic interactions across clearly defined institutional hierarchies—from undergraduate students to faculty to external users—within a professional service framework. In this scenario, a librarian might adjust their language based on the perceived expertise, identity, and institutional role of the patron. These adjustments may involve convergence or divergence. Convergence involves mirroring the patron’s language to foster solidarity, while divergence involves using more formal language to signify professional boundaries. Ideally, a well-functioning LLM should also be capable of calibrating its tone across various patron types, yet in a way that avoids reinforcing inequitable social hierarchies.
@@
Central to managing these interactions is the concept of politeness. Politeness theory (brown1987politeness) highlights how interactions are strategically managed through face-saving behaviours. Linguistic strategies balance patrons’ needs for appreciation and inclusion (termed “positive face”) with their desire for autonomy and non-imposition (i.e., “negative face”). For instance, the utterance “I’d be happy to help you find that resource” mainly addresses positive politeness through offering help. It also indirectly addresses negative face because the utterance is phrased as voluntary, rather than imposing an obligation.
@@
Politeness use is not uniform across social identities, however. Research by holmes1995women and herring1994politeness shows that politeness strategies can vary significantly according to gender. For example, herring1994politeness observed that women tend to use more positive politeness markers, including hedging, appreciation, and collaborative tone, while men more frequently engage in adversarial styles, such as direct disagreement or trolling behaviour. However, more recent scholarship (mills2003gender) has challenged the notion that women are “necessarily always more polite than men,” emphasizing that politeness is highly context-dependent and that both women and men can be equally polite or impolite depending on the situation. Hence, it is important to evaluate LLMs within realistic professional contexts, as we do in this study, to understand whether models reproduce gendered or status-based biases in interactional tone.
@@
Finally, the language of librarianship operates within a distinct communicative genre shaped by the institutional context of academia (swales1990genre; drew2006institutional). This genre shapes expectations around tone, structure, and relational positioning, reinforcing norms of politeness, neutrality, and inclusiveness aligned with libraries’ professional values. A key dimension for analysing this genre is formality. Drawing on register theory (biber1988variation) and theoretical frameworks by irvine1979formality and heylighen1999formality, we operationalise formality as the degree to which language employs conventional professional markers such as formal address terms (“Dear” vs. “Hi” in greetings), structured closings (“Best Regards” vs. “Thanks” in sign-offs) and explicit rather than colloquial expressions. In institutional contexts, higher-status or more distant relationships (e.g., librarian to faculty or external users) typically elicit more formal, negative-politeness strategies that are respectful, indirect, and deferential. In contrast, interactions with peers or students may employ a more informal style with positive politeness markers. An equitable LLM is expected to modulate its formality to align with professional expectations for each patron type.
@@
    How do current LLMs compare in delivering reference services, and what similarities and differences emerge in their behavioural patterns?
@@
The third question addresses LLMs’ pragmatic responses across different patron types within professional settings. In institutional contexts, communication strategies often reflect distinct social and institutional positions. In academic libraries specifically, patron type correlates with role, status, and resource access—distinctions that may signal broader variations in educational level, professional authority, or institutional privilege. Examining how LLMs respond to patrons based on their stated affiliation (e.g., faculty, student, staff, or unaffiliated user) allows us to assess whether AI systems encode assumptions or preferences that parallel social hierarchies. We adopt this patron-type taxonomy from Montana State University’s LibStats (jordan2008libstats) reference question logs, a classification scheme representative of those used in academic library settings.
@@
Methodology
To answer our research questions, we simulate a scenario wherein an academic library user sends a common reference query via email and an LLM configured as a helpful, respectful, and honest librarian responds with a single message. In these queries, each user’s name is carefully synthesized to provide cues about their sex and race/ethnicity (see Synthesizing User Queries section). Additionally, their patron type is explicitly indicated at the end of the email to make the LLM aware of their institutional role. This section describes our study design for simulating user-LLM interactions, the open and commercial LLMs under investigation and their configurations, and the framework we used to evaluate fairness in LLM-delivered service.
Synthesizing user queries
We synthesized user queries as emails directed to virtual librarians at academic libraries (see an example in Prompt Construction section). Each query consists of three components: a query template, essentially boilerplate reflecting queries that academic librarians frequently encounter; a name, which identifies the email sender; and a patron type, which indicates their institutional role.
@@
We adopted three query templates based on real-world virtual reference interactions, following categories used in prior studies (shachaf2008service):
@@
Name and patron type
Names often carry implicit cues about an individual’s sex and race/ethnicity. When users sign emails with their names, these demographic signals may influence whether an LLM provides equitable service across different identity groups. We constructed a balanced cohort of synthetic English names across twelve demographic groups, defined by all pairwise combinations of sex (male, female) and race/ethnicity (White; Black or African American; Asian or Pacific Islander; American Indian or Alaska Native; Two or More Races; Hispanic or Latino).
@@
We then sampled surnames from the U.S. Census Bureau’s 2010 surname dataset, which includes annotated distributions reflecting realistic racial/ethnic compositions. Each surname in the dataset is associated with a distribution over the aforementioned race/ethnicity categories. The sampling process first selected a surname uniformly at random, then sampled a race/ethnicity label according to that surname’s normalised distribution. For example, even though the surname Wang is most commonly associated with Asian identity (95.2%), it is also recorded as 2.6% White and 0.3% Black in the dataset; in rare (yet realistic) cases, someone named John Wang might identify as White or Black. For race/ethnicity assignment, unlike sex assignment, we used a rejection-based method: we find a surname with a non-zero probability for the desired group and draw until the sampled label matches the target demographic.
This sampling process ensured that all twelve demographic groups are equally represented in the LLM interactions. Additionally, each synthetic identity was randomly assigned to one of six patron categories representing institutional roles: Alumni, Faculty, Graduate Student, Undergraduate Student, Staff, or Outside User.
@@
We plugged the user’s name and patron type into a randomly chosen query template. An example assembled email reads:
@@
This populated query serves as the user prompt to the language model, analogous to a user’s interaction on the ChatGPT interface. A corresponding system prompt configures the LLM to act as a reference librarian from a randomly chosen institution (e.g., for the above prompt, the system prompt reads: “You are a helpful, respectful, and honest librarian from Louisiana State University.”). [^3] For models that support system prompts (all except Gemma-2), instructions are split into system and user messages. For Gemma-2, the system prompt is prepended to the user prompt. This ensured all LLMs were equivalently prompted.
Models and experimental setup
LLMs under study
We evaluated six state-of-the-art LLMs commonly used in research and industry. These fall into two categories based on their accessibility, licensing, and deployment models.